The lenght of a rectangular yard is 25ft more than its width. If the width were doubled, the area of the yard would be 1800ft2. Find the dimensions of the yard.
Answer by checkley77(12844)

L=W+25
L*2W=1800
(W+25)2W=1800
2W^2+50W-1800=0
2(W^2+25W-900)=0
2(W-20)(W+45)=0
W-20=0
W=20 FOR THE ORIGINAL WIDTH.
L=20+25=45 FOR THE ORIGINAL LENGTH.
20*2(20+25)=1800
20*2*45)=1800
1800=1800